# Dynamics of contaminant flow through porous media containing biochar adsorbers

### v1

@@ Abstract
Evidence has shown green roof fertilizers released by higher intensity rainfalls contribute as a contaminant to local aquifers. To reduce this, biochar is used as an additive in green roof soil substrates as a fertilizer storage and dispersal regulatory aid. The evolution of contaminant transport and adsorption by biochar added to a packed bed is analyzed using experiments and simulations. Experiment 1 is used to determine the equilibrium capacity and adsorption rate of two types of biochar when immersed in a methylene blue solution. Breakthrough curves of a packed bed of glass beads with randomly interspersed biochar are determined in experiment 2. Simulations are then run to investigate the solute transport and adsorption dynamics at the pore-scale. An analytical model is proposed to capture the behavior of the biochar adsorption capacity and the simulation results are compared with experiment 2. A pore-scale analysis showed that uniformly sized beds are superior in contaminant breakthrough reduction, which is related to the adsorptive surface area and the rate at which adsorption capacity is reached. Cases using the adsorption capacity model display a tight distribution of particle surface concentration at later simulation times, indicating maximum possible adsorption. The beds with dissimilar particle sizes create more channeling effects which reduce adsorptive particle efficiency and consequently higher breakthrough concentration profiles. Comparison between experiments and simulations show good agreement. Improved biochar performance can be achieved by maintaining particle size uniformity alongside high adsorption capacity and adsorption rates appropriate to the rainfall intensity.
@@ 1 Introduction
In the last few decades there has been a renewed interest in the usage of vegetation in urban environments for a variety of reasons, ranging from aesthetic to acoustic to functional. Green roofs are one such vegetated inclusion, consisting of living greenery ranging from full on trees in intensive green roofs to grasses and sedum used in extensive roofs. While there exist inorganic alternatives known as blue or gray roofs, we focus here only the organic alternatives as they require sustenance. The benefits of green roofs in urban environments are well documented, such as their contributions to the reduction in urban noise, air, and water pollution; their effects on the urban heat island as well as building envelopes themselves; and urban rainwater runoff management. These benefits have been covered by a variety of studies over the span of several decades and a more comprehensive review of them may be found in [1].
@@
One of the downsides of using green roofs from a practical standpoint is the need for regular maintenance. This includes potential trimming or more commonly watering and nutrient distribution. Fertilizer is commonly applied to the green roof upon installation, consisting of two variants. A fast-acting fertilizer which gives the plants a boost to survive the initial period of installation and establish themselves; and longer-acting nutrients which aim to keep the additional required maintenance to a minimum. An excess of fertilizer which cannot be stored by the soil or used directly by the vegetation can be carried by rainwater from the soil through the drainage system to locations where it acts as a contaminant rather than a beneficial entity [26]. The balance of sufficient nutrient presence to unwanted contamination requires the soil itself to aid the plants in holding and releasing the applied nutrients while keeping the chemical runoff to a minimum. To this end, biochar has been used to adsorb nutrients, which we call a contaminant or solute, within the soil or growth substrate. Biochar is essentially any organic material that has been carbonized under high temperature through a process known as pyrolysis and can vary in material properties and morphology significantly.
@@
Many variants of biochar have been employed for the purpose of removing contaminants in a soil or packed bed. [5] examined the efficacy of biochar and greenwaste compost on the reduction of zinc, cadmium, and copper within soils and found them to be beneficial in the control of pollutants. [18] provide an early overview on the use of biochar for water and wastewater treatment, comparing adsorption capacities for organic and inorganic contaminants when possible. [2] determined the adsorption of methylene blue (MB) by eucalyptus bark biomass in a packed bed and found that it performs well in the removal of dye-containing effluents. [3] suggested that real contaminants from the mining, pharmaceutical, and agricultural industries should be tested in full column tests where possible. In addition, they reported that pH, surface chemistry, and surface/pore volume all play a role in adsorption capacities of biochar. Each of these playing a more major or minor role dependent upon the organic/inorganic nature of the contaminants. [7] analyzed the adsorption of MB in pine cone biochar in a packed bed column of Kaolin clay and utilized several analytical models to determine breakthrough curves. These curves were compared under varying experimental conditions and the biochar/clay packed bed was found to perform best under low flow rates, high MB concentration, and larger bed depth. [9] determined the efficacy of Capsicum annuum based biochar on the removal of lead from aqueous systems in batch and dynamic flow systems and recommended its use as a cheap alternative for lead removal from fruit juices and wastewater. [27] analyzed competitive sorption and desorption of zinc, cadmium, and lead between compost, biochar, and peat. They found that biochar had the highest adsorption capacity and lowest desorption rate, with lead removal being the most effective. [19] provided a review of the work associated with the use of biomass on the removal of pharmaceutical compounds and reported that biochar made from a variety of plants; such as rice husks, corn, sawdust, and sugar cane, was the most effective in many industrial applications, partly due to their high adsorption capacity.
@@
While the aforementioned studies are valuable in understanding the limits of biochar as an adsorbent for a variety of contaminants, we are unable to fully understand the mechanisms by which they perform best, such as available adsorbing surface area, directing the flow field of the transporting liquid, or their total storage capacity within the micro-porous structure of individual biochar elements. For this level of detail we must utilize numerical models which can describe such systems and provide the information we require. The physical process of contaminant adsorption in a flow field is described mathematically by the advection-diffusion of species, combined with adsorption/desorption; a model which is present across many fields of application. This general model is used to describe processes such as post-combustion CO2 capture [20], modeling the coffee-stain effect during inkjet printing [24], determining optimal geometry of granular heat exchangers [17], analyzing methane adsorption in subsurface shale deposits [14], and describing the behavior of photocatalytic textiles [21] to name a few. Investigations into the dominant factors affecting adsorption rates and capacity have shown that particle and pore morphology play an important role [13, 12]. Adsorption rates themselves also strongly determine the process evolution [15], with lower rates determining adsorption quantity by reaction duration, and higher rates controlling quantity through reaction rate [28]. Non-isothermal systems can result in high heterogeneity of concentration and temperature, with performance suffering from regions of stagnant flow and increased near-wall breakthrough [23]. Many of these studies have focused on the porous medium itself rather than the evolution of the flow and its direct impact on the evolution of a transported species.
@@
When one considers the addition of a reactive species within this system; wherein a solute scalar is adsorbed locally by reactive elements within the porous matrix; understanding the underlying interactions which determine the local concentration distribution is necessary. The mixing itself within porous media has been shown to be complex at high Reynolds numbers (Re). As such, the flow is characterized by an initial advection-dominated regime as the flow penetrates the pore network and is followed by a diffusion-dominated regime wherein the molecular diffusion redistributes the scalar more effectively than the advection. The advective regime consists of stretching of the scalar front in longer finger-like structures, particularly where the flow is strongest, and can be defined by by the plume’s concentration mean, variance, and probability distribution [4]. A thorough overview of current state of research and methods used to solve species transport in porous media can be found in [6].
@@ 2 Experiments
MB, chemical formula C16H18CIN3S, was used as the adsorbate; diluted to 6 mg/L in a phosphate buffer solution containing 380 mg/L KH2PO4 and 300 mg/L K2HPO4. The MB concentration was measured using spectrophotometry with absorbance measured at 680nm. Figure 2 shows the experimental setup that was used for experiments 1 and 2.
@@ 2.1 Experiment 1
Material | Weight [g] | |
Glass | 20.0 | 0.24 |
WCB | 2.0 | 0.42 |
WCB | 3.5 | 0.45 |
WCB | 5.0 | 0.52 |
GB | 2.0 | 0.43 |
GB | 3.5 | 0.47 |
GB | 3.5 | 0.42 |
GB | 5.0 | 0.48 |
GB | 5.0 | 0.49 |
@@
If one wishes to apply a simple capacity model to the system, the resulting boundary condition at the reactive particle surface will be as in equation (4). Note that this condition bears striking similarity to (2) but is distinct, supporting our claim equation (3) is not suitable for solving the mass-transport limited problem. This condition will be formalized in section 3.2 and is only presented here to demonstrate the differences in boundary condition between the experiments.
define quantities and that f(c) is defined later
@@ 3.3 Domain generation and operating conditions
The rainfall intensity of experiment 2 is calculated to be equivalent to 610mm/hr which is clearly far higher than any naturally occurring rainfall event, with the global record sitting around 305mm/hr [25]. Even though such rainfall intensities are higher than any naturally occurring event, it is possible that channeling of runoff to a porous surface could result in such a flow rate, for example. The experimental flow rate results in a Reynolds number (Re) of 0.17 and a Péclet number (Pe) of 250.
Due to the fact that this flow rate far exceeds a realistic rainfall event we reduce these parameters such that Re and Pe, which still allows for a qualitative comparison with the experimental data as Pe , meaning advection and diffusion are of the same order of magnitude. These parameters are equivalent to a rainfall event with intensity 60mm/hr. This condition is used to tune , the diffusivity and viscosity accordingly. It is worth mentioning that common rainfalls are between 3-8mm/hr, however less common extreme events can reach 15 mm/hr or more. We can extrapolate the simulation conditions to lower rainfall intensities by accounting for the time scale of the physical processes involved. The soil will not be as thin as that in the simulations and no runoff infiltration front will be pure contaminant. Additionally, one can alter the diffusivity to compensate such that the physically realistic process is retained.
The resultant mean flow velocity is averaged over the entire domain, given by . The Kozeny-Carman relationship stipulates that if the porosity of a domain is altered the flow rate must be changed accordingly. Since in our cases porosity is approximately unaltered, the resultant mean velocity for the polydisperse case is the same. The biochar percentage is set to 30% in the simulations, double the industry maximum recommendation, to improve statistical analysis on single particle adsorption features without straying significantly from realistic conditions. We match the Damköhler number, , because we wish to have the same balance of flow and reaction rate as in the experiments.
@@
1 | 0.0 | Mono | None |
2 | 0.0 | Poly | None |
3 | 5.3e-6 | Mono | Unlimited |
4 | 5.3e-6 | Poly | Unlimited |
5 | 5.3e-6 | Mono | Limited |
6 | 5.3e-6 | Poly | Limited |
@@ 4.2 Average surface concentration and flux ratio comparison
This hypothesis is given some weight by Figure 8(b) where two experimental cases are directly compared to the simulation cases with limited adsorption capacity. We find a very similar behavior in that both experiments and simulations swiftly reach capacity and subsequently remove little of the solute. Note that there is an initial period of time for the MB solution to reach the porous media through the piping and we do not take wall effects into account in the simulation; both of which are present in the experiment. Consequently, the experimental curves are have a lower gradient and take additional time to reach full capacity, however the steady-state behavior is correctly reflected by the simulations.
@@ 4.5 Particle adsorption capacity
We can now state with reasonable certainty that reactive particles do influence the concentration profiles, particularly the available reactive surface area. An advantage may be achieved by using more mondisperse particles, as they promote a more homogeneous front for the advancing solute and reduce the solute leakage, particularly in the case of particles with limited adsorbing capacity..

### v2

# Dynamics of contaminant flow through porous media containing random adsorbers

@@ Abstract
Many porous media are mixtures of inert and reactive materials, manifesting spatio-chemical heterogeneity. We study the evolution of scalar transport in a chemically heterogeneous material that mimics a green roof soil substrate, fractionally composed of inert and reactive adsorbing particles. These adsorbing particles are equivalent to biochar within a real soil substrate. The scalar transport evolution is determined using experiments and simulations calibrated from experimental data. Experiment 1 is used to determine the equilibrium capacity and adsorption rate of two biochar types when immersed in a methylene blue solution. Breakthrough curves of a packed bed of glass beads with randomly interspersed biochar are determined in experiment 2. Simulations are then run to investigate the solute transport and adsorption dynamics at the pore-scale. An analytical model is proposed to capture the behavior of the biochar adsorption capacity and the simulation results are compared with experiment 2. A pore-scale analysis showed that uniformly sized beds are superior in contaminant breakthrough reduction, which is related to the adsorptive surface area and the rate at which adsorption capacity is reached. Cases using the adsorption capacity model display a tight distribution of particle surface concentration at later simulation times, indicating maximum possible adsorption. The beds with dissimilar particle sizes create more channeling effects which reduce adsorptive particle efficiency and consequently higher breakthrough concentration profiles. Comparison between experiments and simulations show good agreement. Improved biochar performance can be achieved by maintaining particle size uniformity alongside high adsorption capacity and adsorption rates appropriate to the rainfall intensity.
@@ 1 Introduction
Porous media containing a proportion of inert and chemically reactive elements can found in biology in the form of biofilters, bioreactors, and organic tissue. They are also found in geological elements such as rocks or soils, which are also highly chemically heterogeneous, containing a wide variety of chemically dissimilar minerals and/or organic elements. Research into this type of porous media has formerly been primarily driven by the petrochemical industry, however this has shifted in recent years to a more environmental focus. Green roofs are one such example, consisting of living greenery ranging from trees to grasses growing in a soil substrate whose composition can vary widely, from crushed brick to manure to peat moss. The benefits of green roofs in urban environments are well documented, such as their contributions to the reduction in urban noise, air, and water pollution; their effects on the urban heat island as well as building envelopes themselves; and urban rainwater runoff management [1].
@@
Green roofs require regular maintenance, including the administration of fertilizer upon installation and potentially thereafter as the need arises. The fertilizer consists of a fast-acting agent which gives the plants a boost to survive the initial period of installation and establish themselves; and longer-acting nutrients which aim to keep the additional required maintenance to a minimum. Excess fertilizer which cannot be stored by the soil or used directly by the vegetation can be carried by rainwater from the soil through the drainage system to locations where it acts as a contaminant [34]. The addition of biochar to the soil to adsorb excess fertilizer (contaminant or solute) is currently being put forth as viable solution to this problem.
1.1 Biochar in green roofs
Biochar is essentially any organic material that has been carbonized under high temperature through a process known as pyrolysis and can vary in material properties and morphology significantly. A thorough overview of the types of biochar in use commercially can be found in [24].
Biochar is incorporated into green roof soil via three strategies; random application to the top of the bedding material, thorough mixing into the substrate composition prior to vegetation, and set as a layer at the bottom of the substrate. Thorough mixing and layering at the soil base are considered the most effective, with the latter shown to be effective for reducing the leaching of total nitrogen and total phosphorous [17]. The proportion of applied biochar varies, from 5% by weight/volume up to about 40%, however excessive proportions of biochar will have adverse effects on plant growth and contribute to increased contamination [35].
There are few large-scale projects incorporating biochar use on green roofs though a few exist, such as the NWE CASCADE project, in which France intends to implement the use of biochar to enhance stormwater management within the Brittany region. Other similar solutions will likely be implemented within a larger initiative led by Bloomberg Philanthropies as touched upon in [31].
The industrial-scale implementation is dependent upon several factors, such as the ecological impact of large-scale biochar use and the release of additional contaminants present within or on the surface of biochar as a result of its preparation and function. For example, [25] discusses the almost complete removal of ammonia from the soil when excessive biochar is applied, which can be problematic for soil health. Another key factor is availability of biochar itself, given the different feedstock and treatment methods for its creation. A final major consideration is the lifetime of the biochar as an adsorbent and its end-of-life handling. Little work has been done on the effects of aging on biochar performance due to limited long-term projects but what is known is that degrading performance is dependent upon the contaminant and environmental conditions. For example, if the primary goal of the biochar is to trap micro and nano-plastics (MPs and NPs) then both the aging of the biochar and the plastics themselves play an important role in the removal efficiency [15]. Physical degradation of the biochar also occurs during the aging process and leads to biochar dust, which no longer serves its purpose of entrapping pollutants but can act to spread captured plastics or heavy metals due to its increased mobility. It has been shown the granulated biochar in particular is more resistant to this degradation process, thus making it more suitable for applications such as green roofs where the environmental exposure is high [18].
Related to the issue of biochar disposal is the issue of regeneration, which heavily depends upon the captured contaminants. It is possible to regenerate biochar that has captured MPs/NPs using organic solvents such as acetone, however this is not feasible in a green roof environment due to environmental concerns. Water rinsing performs relatively poorly for removing carbon-based adsorbents but can be used [15]. Adsorbed copper can be almost completely recovered [4], as can cadmium [7]. Biochar regeneration over five adsorption-desorption cycles was assessed with regard to volatile organic compounds (VOCs) with reported 88% to 96% regeneration [29]. While these reports are promising, the scalability for some of the processes used for regeneration is in doubt as they are either inappropriate for green roofs due to their open nature, or less efficient if alternatives are used [10].
Biochar may raise the production costs of green roofs, resulting in higher final prices for these products [16]. To evaluate its profitability for both producers and end users, it is essential to compare its effects against using fertilizers alone or in combination with fertilizers [36]. Additionally, given its environmental benefits—such as minimizing nutrient leakage and boosting carbon sequestration—efforts are underway to improve cost calculation models to include both private and societal costs and benefits [6]. However, these calculations are highly case-specific, influenced by factors like location, feedstock, scale, pyrolysis conditions, biochar pricing, and crop type. To promote greater adoption of biochar, [6] suggest developing standardized calculation models and emphasizing additional societal benefits of biochar, such as improved water retention. While the proposed model is comprehensive, it remains unclear how potential hazards associated with biochar [35] will be addressed with it. Thus, the technoeconomic benefits of biochar will be a focus of future research.
1.2 Biochar adsorption capabilities
@@
Many variants of biochar have been employed for the purpose of removing contaminants in a soil or packed bed. [2] determined the adsorption of methylene blue (MB) by eucalyptus bark biomass in a packed bed and found that it performs well in the removal of dye-containing effluents. [9] analyzed the adsorption of MB in pine cone biochar in a packed bed column of Kaolin clay and utilized several analytical models to determine breakthrough curves. These curves were compared under varying experimental conditions and the biochar/clay packed bed was found to perform best under low flow rates, high MB concentration, and larger bed depth. [37] analyzed competitive sorption and desorption of zinc, cadmium, and lead between compost, biochar, and peat. They found that biochar had the highest adsorption capacity and lowest desorption rate, with lead removal being the most effective. [5] examined the efficacy of biochar and greenwaste compost on the reduction of zinc, cadmium, and copper within soils and found them to be beneficial in the control of pollutants. [26] provided a review of the work associated with the use of biomass on the removal of pharmaceutical compounds and reported that biochar made from a variety of plants; such as rice husks, corn, sawdust, and sugar cane, was the most effective in many industrial applications, partly due to their high adsorption capacity. While the aforementioned studies are valuable in understanding biochar as an adsorbent for a variety of contaminants, we are unable to identify the mechanisms by which it performs best, such as available adsorbing surface area, effect on the flow field, or their pore storage capacity. For this level of detail we must utilize numerical models which can describe such systems and provide the level of detail we require.
@@
The physical process of contaminant adsorption in a flow field is described mathematically by the advection-diffusion of species, combined with adsorption/desorption; a model which is present across many fields of application. This general model is used to describe processes such as post-combustion CO2 capture [27], determining optimal geometry of granular heat exchangers [23], analyzing methane adsorption in subsurface shale deposits [20], and describing the behavior of photocatalytic textiles [28] to name a few. Investigations into the dominant factors affecting adsorption rates and capacity have shown that particle and pore morphology play an important role [19, 14]. Adsorption rates themselves also strongly determine the process evolution [21], with lower rates determining adsorption quantity by reaction duration, and higher rates controlling quantity through reaction rate [38]. Non-isothermal systems can result in high heterogeneity of concentration and temperature, with performance suffering from regions of stagnant flow and increased near-wall breakthrough [32].
@@
When one considers a transported reactive species adsorbed locally by reactive elements within a porous matrix; understanding the underlying interactions which determine the local concentration distribution is necessary. The mixing itself within porous media has been shown to be complex at high Reynolds numbers (Re). As such, the flow is characterized by an initial advection-dominated regime as the flow penetrates the pore network and is followed by a diffusion-dominated regime wherein the molecular diffusion redistributes the scalar more effectively than the advection. The advective regime consists of stretching of the scalar front in longer finger-like structures, particularly where the flow is strongest, and can be defined by by the plume’s concentration mean, variance, and probability distribution [3]. A thorough overview of current state of research and methods used to solve species transport in porous media can be found in [8].
@@ 2 Experiments
MB, chemical formula C16H18CIN3S, was used as the adsorbate; diluted to 6 mg/L in a phosphate buffer solution containing 380 mg/L KH2PO4 and 300 mg/L K2HPO4. It is worth noting that due to its properties, methylene blue can be seen as an analog for some organic pollutants however it is by no means representative of all possibilities. Other compounds or the direct pollutant itself may be required for more accuracy and the numerical model must be appropriately modified as well.
The MB concentration was measured using spectrophotometry with absorbance measured at 680nm. Figure 2 shows the experimental setup that was used for experiments 1 and 2.
@@ 2.1 Experiment 1
| Material | Glass | WCB | GB | ||||||
| Weight [g] | 20.0 | 2.0 | 3.5 | 5.0 | 2.0 | 3.5 | 3.5 | 5.0 | 5.0 |
| 0.24 | 0.42 | 0.45 | 0.52 | 0.43 | 0.47 | 0.42 | 0.48 | 0.49 |
@@
If one wishes to apply a simple capacity model to the system, the resulting boundary condition at the reactive particle surface will be as in equation (4). Note that this condition bears striking similarity to (2) but is distinct, supporting our claim equation (3) is not suitable for solving the mass-transport limited problem. In these equations is the molecular diffusivity, Da the Damköhler number, and the capacity model. This condition will be formalized in section 3.2 and is only presented here to demonstrate the differences in boundary condition between the experiments.
@@ 3.3 Domain generation and operating conditions
In order to simulate an intense rainfall event the Péclet number Pe and we assume little inertial influence, thus Re . We set the Damköhler number, , which represents equal magnitudes of molecular diffusion and reaction rates, as calculated in the experiment 2. By applying the conditions required above, the simulation has Re and Pe , which is representative of a rainfall event with intensity of around 60mm/hr. The flow rate of experiment 2 is mL/min, which converts to about 600mm/hr, about double the world record rainfall intensity [33], with Pe and Re . It is worthwhile to note that intensities higher than the record can be reached when runoff is channeled to a porous surface, which is not naturally occurring rainfall. While these values are higher than those chosen for the simulation, the governing physics are identically balanced, as the originally defined conditions are all satisfied, including Da . With the above mentioned flow rates we have been able to measure variations in concentration quantities in a few hours, which we found as a good time lapse to minimize effects from changes in environmental conditions on the measurements.
These conditions are used to tune , the molecular diffusivity and viscosity accordingly. It is worth mentioning that common rainfalls are between 3-8mm/hr, however less common extreme events can reach 15-60 mm/hr or more. We can extrapolate the simulation conditions to lower rainfall intensities by accounting for the time scale of the physical processes involved. The soil will not be as thin as that in the simulations and no runoff infiltration front will be pure contaminant. Additionally, one can alter the diffusivity to compensate such that the physically realistic process is retained.
The resultant mean flow velocity is averaged over the entire domain, given by . The Kozeny-Carman relationship stipulates that if the porosity of a domain is altered the flow rate must be changed accordingly. Since in our cases porosity is approximately unaltered, the resultant mean velocity for the polydisperse case is the same. The biochar percentage is set to 30% in the simulations, double the industry maximum recommendation, to improve statistical analysis on single particle adsorption features without straying significantly from realistic conditions.
@@
| 1 | 0.0 | Mono | None |
| 2 | 0.0 | Poly | None |
| 3 | 5.3e-6 | Mono | Unlimited |
| 4 | 5.3e-6 | Poly | Unlimited |
| 5 | 5.3e-6 | Mono | Limited |
| 6 | 5.3e-6 | Poly | Limited |
@@ 4.2 Average surface concentration and flux ratio comparison
This hypothesis is given some weight by Figure 8(b) where two experimental cases are directly compared to the simulation cases with limited adsorption capacity. We find a very similar behavior in that both experiments and simulations swiftly reach capacity and subsequently remove little of the solute. Note that there is an initial period of time for the MB solution to reach the porous media through the piping and we do not take wall effects into account in the simulation; both of which are present in the experiment. Consequently, the experimental curves have a lower gradient and take additional time to reach full capacity, however the steady-state behavior is correctly reflected by the simulations.
At lower rainfall intensities we expect to see an increased residence time of the adsorbent near the biochar surfaces, however this will not matter if the biochar is already at capacity. In addition to the uncertain reaction kinetics, lower rainfall intensities may also result in unsaturated flow conditions, radically altering the system and introducing capillary forces, making any prediction based upon the current work void. Finally, a lower flow velocity may transition from more complex mixing caused by the microstructure to a more simplified laminar flow in larger pore spaces.
@@ 4.5 Particle adsorption capacity
We can now state with reasonable certainty that reactive particles do influence the concentration profiles, particularly the available reactive surface area. An advantage may be achieved by using more mondisperse particles, as they promote a more homogeneous front for the advancing solute and reduce the solute leakage, particularly in the case of particles with limited adsorbing capacity.
4.6 Modifying the reaction kinetics
We close the paper with a short discussion on the practicalities of accurately modelling the adsorption dynamics present in reality on the surface of biochar. There are several different mechanisms by which adsorption occurs; namely pore-filling, hydrophobic interaction, surface complexation, electrostatic interaction, hydrogen bonding interaction, electron-donor-acceptor (EDA) interaction, and intra-particle diffusion. In addition it has been shown that thermal conditions as well as the pH of the environment play a non-negligible role in the reaction speed [15, 10], as does aging or the presence of dissolved organic matter. The sum total of considering individual biochar characteristics, their intrinsic modes of adsorption, as well as environmental factors renders the modelling a significant challenge. In general current works have made use of first and second order pseudo-kinetic models alongside Langmuir and Freundlich isotherms. Some forays have been undertaken in the realm of applied artificial neural networks and statistical chemical modelling however this is still in its infancy for such an application. If one wishes to enhance the models governing the reaction kinetics we need more experimental data to determine which of the above modes of adsorption should be considered and how the environmental and aging aspects can be included in the models. While some kind of pre-factors can be easily applied they are not representative of any physical phenomenon and some possible empirical model msut be constructed which accurately captures these effects may be a simple and efficient solution. Alternatively the full chemical species and reactions can be included in the modeling but this undertaking may not be compatible with how the lattice Boltzmann method operates but would ensure a more chemically consistent interaction. Pore filling and intra-particle diffusion might be handled via some volumetric functions that account for internal pore volume and pore network but this is also impractical to implement. Whichever strategy is used, more experimental data is likely to be required and understanding the fundamental physical interactions in such a system are a good place to aid in determining where to next add the increased realism required for more accurate and applicable predictions.